Solve the following word problem. Be sure to show the equation you use for the solution and your work.
Geometry. The length of a rectangular playing field is 5 ft less than twice its width. If the perimeter of the playing field is 230 ft, find the length and width of the field.
This is what I have so far.
2*w -5ft=L
P=230ft

Answer by checkley75(3666)

L=2W-5
2L*2W=230
2(2W-5)+2W=230
4W-10+2W=230
6W=230+10
6W=240
W=240/6
W=40 FT IS THE WIDTH
L=2*40-5
L=80-5
W=75 FT IS THE LENGTH
PROOF
2*40=2*75=230
80+150=230
230=230